# Convergence in charmonium structure: light-front wave functions from basis light-front quantization and Dyson-Schwinger equations

### v1

@@ I Light-front wave functions
$\psi_{s\bar{s}/P}(x,\vec{k}_{\perp})=\frac{1}{2p^{+}}\int{\mathrm{d}}^{3}ze^{{\frac{i}{2}}xp^{+}z^{-}-i\vec{k}_{\perp}\cdot\vec{z}_{\perp}}\\ \langle 0|\bar{\psi}(+{\frac{z}{2}})\gamma^{+}\gamma_{5}\psi(-{\frac{z}{2}})|P(p)\rangle_{z^{+}=0},$ (4)
@@ II Form factors and hadronic densities
[…]
@@
where, $\tensor{\nabla}_{\perp}$ only acts on the suppressed wave functions and $f\tensor{\nabla}g=f\nabla g-(\nabla f)g$.
Fig. 3 compares the gravitational form factors $A(Q^{2})$ and $D(Q^{2})$ computed using LFWFs from BLFQ and from DSE. As one can see, the BLFQ and DSE results are in good agreement. The $D$-term, defined as $D\equiv D(0)$, is known to be sensitive to the interaction. In Ref. [83], the authors argue that the $D$-term of the pseudoscalar meson should lie within $(-1,-1/3)$, which are the values in the chiral limit and the infinite mass limit respectively. Using DSE with a contact interaction, they obtain $D=-0.58$ for $\eta_{c}$. In contrast, the values extracted from BLFQ and from DSE LFWFs in this work are $D_{\textsc{blfq}}=-4.6$ and $D_{\textsc{dse}}=-4.1$.
@@ III Light-cone distribution amplitudes and radiative transitions
$\langle 0|\bar{\psi}(-{\frac{1}{2}}z)\gamma^{+}\gamma_{5}\psi(+{\frac{1}{2}}z)|P(p)\rangle_{\mu}\\ =if_{P}p^{+}\int_{0}^{1}{\mathrm{d}}xe^{{\frac{i}{2}}xp^{+}z^{-}}\phi_{P}(x,\mu).$ (16)
@@
Figure 4: (Colors online) Comparison of the leading-twist LCDA of $\eta_{c}$ as obtained from BLFQ ($N_{\text{max}}=8$, solid lines) and from DSE (dashed lines). The uncertainty bands are computed from the difference between $N_{\text{max}}=8$ and $N_{\text{max}}=16$ results in BLFQ. The dot-dashed lines correspond to the pQCD asymptotic form $6x(1-x)$ [7]. The DSE results are evolved from $\mu_{0}=2.6\,\mathrm{GeV}$ to $\mu=2.8\,\mathrm{GeV}$ to match the BLFQ energy scale.

### v2

# thanks: Corresponding author

@@ I Light-front wave functions
$\psi_{\uparrow\downarrow-\downarrow\uparrow/P}(x,\vec{k}_{\perp})=\sqrt{\frac{x(1-x)}{2}}\int{\mathrm{d}}^{3}ze^{{\frac{i}{2}}xp^{+}z^{-}-i\vec{k}_{\perp}\cdot\vec{z}_{\perp}}\\ \langle 0|\frac{1}{\sqrt{N_{c}}}\bar{\psi}(0)\gamma^{+}\gamma_{5}\psi(z)|P(p)\rangle_{z^{+}=0,\vec{p}_{\perp}=0},$ (4)
@@ II Form factors and hadronic densities
where, the $\langle\cdots\rangle$ represents the quantum average:
@@
[…]
Fig. 3 compares the gravitational form factors $A(Q^{2})$ and $D(Q^{2})$ computed using LFWFs from BLFQ and from DSE. As one can see, the BLFQ and DSE results are in good agreement. The $D$-term, defined as $D\equiv D(0)$, is known to be sensitive to the interaction. In Ref. [83], the authors argue that the $D$-term of the pseudoscalar meson should lie within (1,1/3)(-1,-1/3), which are the values in the chiral limit and the infinite mass limit respectively. Using DSE with a contact interaction, they obtain D=0.58D=-0.58 for $\eta_{c}$. In contrast, the values extracted from BLFQ and from DSE LFWFs in this work are $D_{\textsc{blfq}}=-4.6$ and $D_{\textsc{dse}}=-4.1$.
@@ III Light-cone distribution amplitudes and radiative transitions
$\langle 0|\bar{\psi}(+{\frac{1}{2}}z)\gamma^{+}\gamma_{5}\psi(-{\frac{1}{2}}z)|P(p)\rangle_{z^{+},\vec{z}_{\perp}=0,\mu}\\ =if_{P}p^{+}\int_{0}^{1}{\mathrm{d}}xe^{{\frac{i}{2}}(x-\frac{1}{2})p^{+}z^{-}}\phi_{P}(x,\mu).$ (16)
@@
Figure 4: (Colors online) Comparison of the leading-twist LCDA of $\eta_{c}$ as obtained from BLFQ ($N_{\text{max}}=8$, solid lines) and from DSE (dashed lines). The uncertainty bands are computed from the difference between $N_{\text{max}}=8$ and $N_{\text{max}}=16$ results in BLFQ. The dot-dashed lines correspond to the pQCD asymptotic form 6x(1x)6x(1-x) [7]. The DSE results are evolved from $\mu_{0}=2.6\,\mathrm{GeV}$ to $\mu=2.8\,\mathrm{GeV}$ to match the BLFQ energy scale.